OpenStudy (loser66):

Find real and imaginary part of \((\dfrac{-1-i\sqrt3}{2})^6\) by polar expression. Please, help

OpenStudy (loser66):

oh, I know what is wrong, it should be \(z = |-1| e^{i4\pi/3}\) right?

OpenStudy (irishboy123):

\((\dfrac{-1-i\sqrt3}{2})^6\)

OpenStudy (loser66):

Yes

OpenStudy (anonymous):

idk why i typed this :O

OpenStudy (anonymous):

and @Loser66 the polar form is the one with cosine and sine not exponential Form

OpenStudy (loser66):

But we can find the real part and imaginary part by polar, right?

OpenStudy (loser66):

As above, I have \(z = e^{4i\pi/3}\) , then \(z^6 = e^{i8\pi}= cos (8\pi) + isin(8\pi)\) Hence R(z) =1 and Im (z) =0, right?

OpenStudy (anonymous):

yeah sure \( (x+iy)^n=(r\cos n \theta+i r \sin n \theta) \)

OpenStudy (irishboy123):

\[(\dfrac{-1-i\sqrt3}{2})^6\] \[= (-\frac{1}{2} -i\frac{\sqrt{3}}{2})^6\] \[=(e^{i \frac{4 \pi}{3}})^6\] \[=e^{i 8 \pi} = 1\]

OpenStudy (irishboy123):

find the 6 sixth roots of 1, next that's more frustrating.

OpenStudy (irishboy123):